A box with a total surface area of 0.83 m2
and
a wall thickness of 2.72 cm is made of an insulating material. A 11.8 W electric heater
inside the box maintains the inside temperature at 7.4
◦C above the outside temperature.
Find the thermal conductivity of the insulating material.
Answer in units of W/m · ◦ C.
1
Expert's answer
2020-04-09T09:27:24-0400
"k=\frac{qL}{A\Delta T}"

"k=\frac{(11.8)(0.0272)}{(0.83)(7.4)}=0.052\frac{W}{m\degree C}"